sophiaespinosa:

△ABC is dilated from the origin by scale factor r=1/4. Draw △A'B'C' and label the vertices with coordinates. the images triangle is A=(-10,6) B=(-11,2) C=(-4,4)

snowflake0531:

for each coordinate (x,y) multiply the x and y by 1/4 (x,y) -> (1/4 x , 1/4 y)

sophiaespinosa:

mhm

snowflake0531:

do you understand?

sophiaespinosa:

yes i understan that i have to do that but i forgiot hot to multiply it my 1/4

snowflake0531:

A = (-10,6) \( A=(-10 * \frac{1}{4} , 6 * \frac{1}{4}) = (\frac{-10}{4},\frac{6}{4})\) divide numerator and denominator by 2 \[(\frac{-5}{2},\frac{3}{2})\]

snowflake0531:

do the same with the other coordinates

sophiaespinosa:

oh ok thank you so much^_^

snowflake0531:

yw~

sophiaespinosa:

^_^^-^ thank you I really appreciate it

sophiaespinosa:

sorry to disturb so i divide everything by 2 or are each of them different

darkknight:

B = (-11, 2), scale factor of 1/4 (x,y) -> (1/4 x , 1/4 y) so (1/4 * -11, 1/4 * 2) = (-11/4 , 2/4) can you simplify the x or the y value?

sophiaespinosa:

that is what i am confused about I dont know how to do that

sophiaespinosa:

I was thinking -7,2 but i dont know if it is right